OpenStudy (anonymous):

Differentiate: y=1+cosx/1-cosx

myininaya (myininaya):

\[(\frac{1+\cos(x)}{1-\cos(x))})'=\frac{(1+\cos(x))'(1-\cos(x))-(1+\cos(x))(1-\cos(x))'}{(1-\cos(x))^2}\]

OpenStudy (anonymous):

1+cosx/1-cosx = cot( x/2) now diffnciate it...

OpenStudy (anonymous):

sorry cot^2(x/2)

myininaya (myininaya):

\[(1+\cos(x))'=0-\sin(x) =-\sin(x)\] \[(1-\cos(x))'=0-(-\sin(x))=+\sin(x)=\sin(x)\] And I suppose we could had attempted simplifying before differentiating lol

OpenStudy (anonymous):

\[-2\sin \chi \div \left( 1-\cos \chi \right)^{2}\]

OpenStudy (anonymous):

??

myininaya (myininaya):

\[\frac{-\sin(x)(1-\cos(x))-(1+\cos(x))(\sin(x))}{(1-\cos(x))^2}=\frac{\sin(x)(-(1-\cos(x)-(1+\cos(x))}{(1-\cos(x))^2}\] \[=\frac{\sin(x)(-2)}{(1-\cos(x))^2}\] yes thats right! :)